OpenStudy (anonymous):

an internet search engine search for a certain key word in a sequence of independent web sites . it is believed that 20% of the web sites contain this keyword a) compute the probability that atleast 5 of the first 10 sites contain the given key word b) compute the probability that the search engine had to visit at least 5 sites in order to find the first occurrence of a keyword?

OpenStudy (anonymous):

Have you tried solving it first?

OpenStudy (anonymous):

yeah but cant understand it plz help

OpenStudy (anonymous):

a.)1/2 b.)1/5 generally, probability of outcomes of E is just P(E)= total no. of outcomes from e/total number of outcomes I think

OpenStudy (anonymous):

@annas hi bro.

OpenStudy (anonymous):

hay safi wassup!

OpenStudy (ash2326):

@annas do you have the answers for this, I think what @anonymous has posted aren't correct

OpenStudy (anonymous):

yeah they are not correct 0.032 and 0.40

OpenStudy (ash2326):

It's given that 20 % of the websites have the keyword So Probability of finding the keyword on a website is \[P=20\%=\frac{20}{100}=\frac{1} {5}\] Probability of not finding is \[1-\frac{1}{5}=\frac{4}{5}\]

OpenStudy (ash2326):

a) compute the probability that atleast 5 of the first 10 sites contain the given key word So we have to find the probability of finding keyword on 5, 6, 7, 8, 9 and 10 sites P (5) is the probability of finding keyword on 5 sites and not finding on 10 \[P(5)=(\frac{1}{5} )^5(\frac{4}{5})^5\] similarly \[P(6)=(\frac{1}{5} )^6(\frac{4}{5})^4\] \[P(6)=(\frac{1}{5} )^7(\frac{4}{5})^3\]\[P(6)=(\frac{1}{5} )^8(\frac{4}{5})^2\]\[P(6)=(\frac{1}{5} )^9(\frac{4}{5})^1\]\[P(6)=(\frac{1}{5} )^{10}\] So total probability=P(5)+P(6)+P(7)+P(8)+P(9)+P(10)

OpenStudy (anonymous):

thanks @ash2326 really own you one for helping

OpenStudy (ash2326):

But check the answer, I'm not getting 0.032

OpenStudy (anonymous):

i'll check it. by the way how much you r getting

OpenStudy (ash2326):

0.000139

OpenStudy (anonymous):

then i think we must use some distribution

OpenStudy (ash2326):

b) compute the probability that the search engine had to visit at least 5 sites in order to find the first occurrence of a keyword? P=1-(P(1)+P(2)+P(3)+P(4)) P(1)= probability of finding on first website= \(\frac{1}{5}\) P(2)=probability of finding on second website=\(\frac{4}{5} \times \frac{1}{5}\) P(3)=probability of finding on third website=\(\frac{4}{5} \times \frac{4 }{5} \times \frac{1}{5}\) P(4)=probability of finding on fourth website=\(\frac{4}{5} \times \frac{4 }{5} \times \frac{4 }{5} \times \frac{1}{5}\) We have \[P=1-(\frac{1}{5}+\frac{4}{5} \times \frac{1}{5}+\frac{4}{5} \times \frac{4}{5}\times \frac{1}{5}+\frac{4}{5} \times \frac{4}{5}\times \frac{4}{5} \times \frac{1}{5})\]

OpenStudy (anonymous):

@ash nice buddy good going @annas in the first part atleast word is used so you have to use binomail distribution ncx p^x(q)^n-x

OpenStudy (anonymous):

yeah i know that @deadman340